A spring is used to fire a 100-g block horizontally. The spring has a spring constant of 20 N/m and is initially compressed by 5.0 cm. The block is initially at rest.

a) What is the initial elastic potential energy?

b) The spring is released. The block leaves the spring when the spring is at its relaxed state. What is the kinetic energy of the block as it leaves the spring?

c) After the block leaves the spring, it enter a surface. The surface has a coefficient of kinetic friction of 0.2. How far will the block travel on this surface before it will stop?
Answer #1

a)

U = 0.5 k x^2 = 0.5* 20* 0.05^2 = 0.025 J

b)

kE = spring potential energy = 0.025 J

c)

using work-energy theorem

u mg d = kE

0.2* 0.1* 9.8* d = 0.025

0.196 d = 0.025

d = 0.025 / 0.196

d = 0.12755 m

======

Comment before rate in case any doubt, will reply for sure.. goodluck